OpenStudy (anonymous):

Ted Thomason found a good deal when he purchased a bedding set. He paid the store's clerk two $20 bills and a $5 bill. The clerk gave him $2.65 in change. If the sales tax is 7.2%, what was the selling price of the bedding set before taxes?

OpenStudy (anonymous):

i guess he paid a total of \(2\times 20+5-2.65=45-2.65=\$42.35\)

OpenStudy (anonymous):

to add \(7.2\%\) you multiply a number by \(1.072\) and so you know \[1.072P=42.35\] making \[P=42.35\div 1.072\]

OpenStudy (anonymous):

So $39.51

OpenStudy (anonymous):

or \(\$39.50\) rounding down slightly either way, you are probably more correct